A wheel of radius a, with its mass concentrated on

?explanations needed, thanks.
Answer #1

using equations from centripetal and centrifugal forces are

entripetal force = mv2 / R

centrifugal force = mv2 / R

and the total force is F

F = mv2 / R + mv2 / R

F = 2mv2 / R ---- 1

and the force is with an angle \alpha

F = m g sin\alpha / cos\alpha

F = m g tan\alpha ----- 2

equating equations 1 and 2

m g tan\alpha = 2mv2 / R

g tan\alpha = 2 v2 / R

R = 2 v2 / g tan\alpha

here given \alpha = 300 , v = 5 m/s

and g = 9.8 m/s2

then substituting values in the above equation

R = 2 X 5 X 5 / 9.8 X tan30

R = 50 / 9.8 X 0.577

R = 50 / 5.658

R = 8.837 m